OpenStudy (anonymous):

how would you solve the problem f(x)= x-2/x when f(x)=-1?

OpenStudy (anonymous):

could someone assist

OpenStudy (anonymous):

f(x)= x-2/x so when f(x)=-1 -> -1 = x-2/x -x = x-2 -2x=-2 -> x=1